# Robust Transmission Design for RIS-Assisted Integrated Sensing and Communication Systems

### v1

@@ III-C Transmit Beamforming
The RIS beamforming matrix is fixed in this subsection. The transmit beamforming problem is formulated as a standard semidefinite programming (SDP) problem, which can be solved by the SDR method. Specifically, the transmit beamforming problem can be written as
@@
where $\boldsymbol{\Gamma}_{k}=\boldsymbol{s}_{k}\boldsymbol{s}_{k}^{H}$. The problem in (32a) can be transformed to a SDP problem by omitting the (32d) constraint. Then, the SDP problem can be solved by the CVX tool using the inter point method (IPM). We use rank-1 approximation method, such as the Gaussian randomization, to obtain the optimal $\boldsymbol{s}_{k}^{*}$ after we obtain the optimal $\boldsymbol{\Gamma}_{k}^{*}$.
@@ III-D1 Scheme 1
[…]
@@ III-D2 Scheme 2
In this scheme, the RIS beamforming problem is transformed into a square penalty (SP) problem in order to accelerate the convergence of the AO algorithm. The SP problem is expressed as
@@
where $\tilde{\mathcal{F}}$ is the convex hull of $\mathcal{F}$, i.e., $\tilde{\mathcal{F}}$ is a regular polygon with $\mathcal{F}$ as its vertices. Note that the phase shifts for RIS elements will be located at the vertices of the regular polygon when the penalty factor $\lambda$ is sufficiently large. The objective function in (40a) is formulated from (31b) and (31c), where the constant terms, such as $\text{Tr}(\boldsymbol{\Psi}_{k})$ and $y_{k}$, are omitted. The basic idea of (40a) is to minimize the outage probability of all communication users. Then, we use gradient extrapolated majorization-maximization (GEMM) method to solve the SP problem. The GEMM method is an efficient algorithm based on gradient ascent to solve the smooth but non-convex problem [46].
@@
where $i$ denotes the iterative number. We use the accelerated projected gradient (APG) method to solve (III-D2), since the APG method has better convergence performance than the projected gradient (PG).
@@
where $\Pi_{\tilde{\mathcal{F}}}$ is the projection of $\boldsymbol{\theta}$ onto $\tilde{\mathcal{F}}$, $\beta^{(i)}$ denotes the step length, and $\boldsymbol{z}^{(i)}$ denotes the extrapolated point of the majorant function at the $i$-th iteration. It should be noted that a large number of iterations are needed to solve (III-D2). Considering the computational complexity, we only utilize one step APG method to update (43). Moreover, $\boldsymbol{z}^{(i)}$ is computed by
@@
The proposed algorithm for solving the RIS beamforming problem (III-D2) based on GEMM method is presented in Algorithm 1. The proposed AO algorithm for solving the robust beamforming problem (III-A) is presented in Algorithm 2.
@@
while $i\leq i_{\text{max}}$, and (III-D2) is satisfied do
@@
    7. Obtain the gradient $\nabla_{\boldsymbol{\theta}}F_{\lambda}$ using (III-D2).
    8. Choose the step length $\beta^{(i)}$ by (III-D2).
    9. Obtain $\boldsymbol{\theta}^{(i+1)}$ by (43) and (III-D2).
@@ IV-A Complexity Analysis for SDP Problems
The SDP problems (32a) and (38a) can be solved with a worst-case complexity of $\mathcal{O}_{\text{SDP}}^{\text{Transmit}}=\mathcal{O}\left(N^{4}K^{0.5}+K^{4.5}\right)$ and $\mathcal{O}_{\text{SDP}}^{\text{RIS}}=\mathcal{O}\left(M^{4}K^{0.5}+K^{4.5}\right)$, respectively.
@@ IV-B Complexity Analysis for the SP Problem
We will show the convergence performance for the proposed AO algorithm by simulation results in the next section.
@@ V Simulation Results
We assume that the BS is located at(0m, 0m); the RIS is located at (50m,10m); the $K$ users and $L$ targets are uniformly distributed within a circle centered at (70m,0m) with a radius of 5 meters and within a circle centered at the base station with a radius of 70 meters. Unless otherwise specified, the number of DFRC BS antennas is 16, the number of targets is 2, the number of users is 2, the communication rate threshold $R_{0}$ equals 2 bits/s/Hz, the CRB threshold equals 0.01, the probability $\rho_{k}$ and $p_{l}$ both equal 0.05, and the noise power is -110dBm.
@@
The trends of all curves in Fig. 2 are consistent. When the number of RIS elements is small (less than about 10), the transmit power decreases as the number of RIS elements increases. However, when the number of RIS elements is larger than about 10, the transmit power increases with the number of RIS elements. This is because when the number of RIS elements is small, the transmit power used to compensate for CSI error is also small. Nevertheless, when the number of RIS elements is large, the power loss caused by compensating for CSI error is greater than the power gain provided by the RIS, resulting in an increase of the transmit power as the number of RIS elements increases. We will explain later why the transmit power decreases when the number of RIS elements is greater than about 50. Furthermore, the transmit power of the RIS continuous phase shifts scheme in Fig. 2 is lower than that of RIS beamforming scheme 1 and RIS beamforming scheme 2, as expected. The results in Fig. 2 break the intuitive conclusion that the more RIS elements, the greater the gain. When perfect CSI cannot be obtained, we should carefully adjust the number of RIS elements.
@@
Fig. 5 depicts the performance of the AO algorithms based on RIS continuous phase shifts, RIS beamforming scheme 1, and RIS beamforming scheme 2. All schemes achieve convergence within 20 AO iterations. Notably, the AO algorithm based on RIS beamforming scheme 1 achieves convergence in 15 iterations, while the AO algorithm based on RIS beamforming scheme 2 achieves convergence in only four iterations.
Figure 6 illustrates that all schemes achieve 100% feasibility when the number of RIS elements is below 34. However, as the number of RIS elements increases beyond 34, the feasibility of all schemes rapidly decreases. Only the channel realizations with good channel conditions are feasible when the number of RIS elements is large. The good channel conditions reduce the required transmit power, which is why the transmit power decreases as the number of RIS elements increases in Fig. 2 and 3.
@@ VI Conclusion
We propose a decomposition-based large deviation inequality approach to approximate non-convex outage probability constraints for communication and non-convex probability constraints for the sensing CRB. An AO algorithm is then employed to solve the joint beamforming problem. Specifically, we introduce an SP-based GEMM algorithm to solve the discrete RIS beamforming problem. Simulation results demonstrate the excellent performance of the proposed GEMM-based AO algorithm. It should be noted that increasing the number of RIS elements does not lead to better performance when the perfect CSI is not available. Thus, we should carefully select the number of RIS elements to guarantee communication rate and CRB for sensing parameter estimation. In our future work, we aim to address the infeasibility issue of the proposed algorithm for a large number of RIS elements. This can be accomplished by employing cognitive radio technology to dynamically adjust communication and sensing parameters, such as bandwidth and modulation order, as well as the number of RIS elements, to adapt to the imperfect CSI. This paper presents a novel approach for designing robust transmission schemes in RIS-assisted ISAC systems.
@@
The second part of (III-B1) is rewritten as
@@
where we user the property of the trace-operator and the vec-operator [47, Eq. (521)], i.e., $\boldsymbol{a}^{T}\boldsymbol{Bc}=\text{Tr}(\boldsymbol{ac}^{T}\boldsymbol{B})=\text{vec}(\boldsymbol{ca}^{T})^{T}\text{vec}(\boldsymbol{B})$. We can formulate (Appendix B Proof of Proposition 1) as $2\text{Re}(\boldsymbol{t}_{k}^{H}\boldsymbol{e}_{k})$ using the definition in (19) and (III-B1).
The third part of (III-B1) is rewritten as
@@
where we use the property of the vec-operator [47, Eq. (524)], i.e., $\boldsymbol{a}^{T}\boldsymbol{XBX}^{T}\boldsymbol{c}=\text{vec}(\boldsymbol{X})^{T}\left(\boldsymbol{B}\otimes\boldsymbol{ca}^{T}\right)\text{vec}(\boldsymbol{X})$. Please refer to B.1.1 of [47] for a proof for this property. We can formulate (Appendix B Proof of Proposition 1) as $\boldsymbol{e}_{k}^{H}\boldsymbol{Q}_{k}\boldsymbol{e}_{k}$ using the definition in (19) and (20).
The first part and the last part of (III-B1) form $s_{k}$. Thus, we obtain (III-B1). This completes the proof of Proposition 1.
@@
In order to derive a convex restriction of (III-B1) using Lemma 1, we set $\boldsymbol{x}=\boldsymbol{e}_{k}$, $\boldsymbol{Q}=\boldsymbol{Q}_{k}$, $\boldsymbol{r}=\boldsymbol{r}_{k}$, $\eta_{k}=\text{Tr}(\boldsymbol{Q}_{k})+s_{k}$, and $T_{k}=v_{k}\|\boldsymbol{Q}_{k}\|_{F}+\frac{1}{\sqrt{2}}\|\boldsymbol{r}_{k}\|$. We choose $v_{k}$ as the solution to $[1-1/(2v_{k}^{2})]v_{k}=\sqrt{\ln(1/\rho_{k})}$. We could observe that when $\eta_{k}\textgreater 2\bar{\theta}_{k}v_{k}T_{k}$, i.e., $\eta_{k}\geq 2\sqrt{\ln(1/\rho_{k})}T_{k}$, the outage probability constraint (III-B1) is always satisfied. It is shown as the following equation.
@@
Further, (III-B1) is also satisfied if $\eta_{k}=2\bar{\theta}_{k}v_{k}T_{k}$. However, (III-B1) may not be satisfied if $\eta_{k}\textless 2\bar{\theta}_{k}v_{k}T_{k}$. Thus, we add constraint for this situation to ensure the outage probability is no more that $\rho_{k}$. The constraint is expressed as

### v2

@@ III-C Transmit Beamforming
The RIS beamforming matrix is fixed in this subsection. The transmit beamforming problem is formulated as a standard SDP problem, which can be solved by the SDR method. Specifically, the transmit beamforming problem can be written as
@@
where $\boldsymbol{\Gamma}_{k}=\boldsymbol{s}_{k}\boldsymbol{s}_{k}^{H}$. The problem in (32a) can be transformed to a SDP problem by omitting the (32d) constraint. Then, the SDP problem can be solved by the CVX tool using the inter point method (IPM). By checking the Karush-Kuhn-Tucker (KKT) conditions of the problem in (32a) without the rank one constraints, it is shown that the optimal beamforming matrices, i.e., $\boldsymbol{\Gamma}_{k}^{*},\forall k\in\mathcal{K}$, satisfying $\text{rank}(\boldsymbol{\Gamma}_{k}^{*})=1$ can always be obtained. This means we can always find the optimal BS beamforming matrix through the SDR method.
@@ III-D1 Scheme 1
$\bar{\boldsymbol{\theta}}=\begin{bmatrix}\boldsymbol{\theta}\\ t\end{bmatrix},\bar{\boldsymbol{G}}_{k}=\begin{bmatrix}\hat{\boldsymbol{H}}_{\text{BRU},k}\boldsymbol{\Psi}_{k}\hat{\boldsymbol{H}}_{\text{BRU},k}^{H},&\!\!\!\!\!\hat{\boldsymbol{H}}_{\text{BRU},k}\boldsymbol{\Psi}_{k}\hat{\boldsymbol{h}}_{\text{BU},k}\\ \hat{\boldsymbol{h}}_{\text{BU},k}^{H}\boldsymbol{\Psi}_{k}\hat{\boldsymbol{H}}_{\text{BRU},k}^{H},&\!\!\!\!\!0\end{bmatrix},$ (35)
@@
Since the SDP problem is not guaranteed to obtain a beamforming matrix with rank on, a Gaussian randomization algorithm is proposed to ensure a rank one solution. We present the Gaussian randomization algorithm in Algorithm 1.
@@
Algorithm 1 Gaussian Randomization Method.
  1. Inputs: $\boldsymbol{\bar{\Theta}}$,$\hat{\boldsymbol{h}}_{\text{BU},k}$, $\hat{\boldsymbol{H}}_{\text{BRU},k}$,$\hat{\alpha}_{l}$, $\gamma_{\text{BU},k}^{2}$, $\gamma_{\text{BRU},k}^{2}$, $\varepsilon_{l}^{2}$, $r_{k}$, $\rho_{k}$, $p_{l}$, $\phi_{l}$, $\sigma_{\text{com},k}^{2}$, $\sigma_{\text{sen}}^{2}$, $G_{\text{max}}$.
  2. Outputs: Beamforming matrix $\boldsymbol{\Theta}^{*}$ for (33a).
  3. Eignvalue decompose $\boldsymbol{\bar{\Theta}}$, i.e., $\boldsymbol{\bar{\Theta}}=\boldsymbol{\Sigma}\boldsymbol{V}$, where $\boldsymbol{\Sigma}$ is the eignvalue matrix and $\boldsymbol{V}$ is the eigenvector matrix.
  for $i=1$ to $G_{\text{max}}$ do
     4. Randomly generate $\tilde{\boldsymbol{x}}_{i}$, which is assume to follow the standard CSCG distribution. And let $\tilde{\boldsymbol{\theta}}_{i}=\boldsymbol{V}\boldsymbol{\Sigma}^{1/2}\tilde{\boldsymbol{x}}_{i}$.
     if $\tilde{\boldsymbol{\theta}}_{i}$ makes the communication rate bigger then
        5. Let $\boldsymbol{\Theta}^{*}=\text{diag}\left(\tilde{\boldsymbol{\theta}}_{i}\right)$.
     end if
  end for
@@ III-D2 Scheme 2
In this scheme, the RIS beamforming problem is transformed into a SP problem in order to accelerate the convergence of the AO algorithm. The SP problem is expressed as
@@
where $\tilde{\mathcal{F}}$ is the convex hull of $\mathcal{F}$, i.e., $\tilde{\mathcal{F}}$ is a regular polygon with $\mathcal{F}$ as its vertices. Note that the phase shifts for RIS elements will be located at the vertices of the regular polygon when the penalty factor $\lambda$ is sufficiently large. The objective function in (40a) is formulated from (31b) and (31c), where the constant terms, such as $\text{Tr}(\boldsymbol{\Psi}_{k})$ and $y_{k}$, are omitted. The basic idea of (40a) is to minimize the outage probability of all communication users. Then, we use GEMM method to solve the SP problem. The GEMM method is an efficient algorithm based on gradient ascent to solve the smooth but non-convex problem [48].
@@
where $i$ denotes the iterative number. We use the APG method to solve (42), since the APG method has better convergence performance than the projected gradient (PG).
@@
where $\Pi_{\tilde{\mathcal{F}}}$ is the projection of $\boldsymbol{\theta}$ onto $\tilde{\mathcal{F}}$, $\beta^{(i)}$ denotes the step length, and $\boldsymbol{z}^{(i)}$ denotes the extrapolated point of the majorant function at the $i$-th iteration. It should be noted that a large number of iterations are needed to solve (42). Considering the computational complexity, we only utilize one step APG method to update (43). Moreover, $\boldsymbol{z}^{(i)}$ is computed by
@@
The proposed algorithm for solving the RIS beamforming problem (40) based on GEMM method is presented in Algorithm 2. The proposed AO algorithm for solving the robust beamforming problem (14) is presented in Algorithm 3.
@@
  while $i\leq i_{\text{max}}$, and (48) is satisfied do
@@
     7. Obtain the gradient $\nabla_{\boldsymbol{\theta}}F_{\lambda}$ using (49).
     8. Choose the step length $\beta^{(i)}$ by (48).
     9. Obtain $\boldsymbol{\theta}^{(i+1)}$ by (43) and (47).
@@ IV-A Complexity Analysis for SDP Problems
The SDP problems with $n$ constraints and the dimension of variables in each constraint equals to $m\times m$ can be solved with a worst-case complexity of $\mathcal{O}\left(\log\left(1/\eta\right)\max\left\{m,n\right\}^{4}m^{0.5}\right)$, where $\eta$ is the accuracy of solution [50]. Therefore, the SDP problems (32a) and (38a) can be solved with a worst-case complexity of $\mathcal{O}_{\text{SDP}}^{\text{Transmit}}=\mathcal{O}\left(\log\left(1/\eta\right)\max\left\{N,4K+L\right\}^{4}N^{0.5}\right)$ and $\mathcal{O}_{\text{SDP}}^{\text{RIS}}=\mathcal{O}\left(\log\left(1/\eta\right)\max\left\{M+1,3K+2\right\}^{4}(M+1)^{0.5}\right)$, respectively.
@@ IV-B Complexity Analysis for the SP Problem
IV-C Convergence Analysis for the AO-SDR Algorithm
In this subsection, we analyze the convergence performance of the proposed AO-SDR algorithm, i.e., the AO based on the SDR method and the RIS beamforming scheme 1. Firstly, the transmit power obtained by the proposed AO-SDR algorithm has a lower bound due to user rate constraints. We assume the lower bound of the transmit power to be $P_{\text{min}}$. Next, we prove that with the iterations of the AO algorithm, the transmit power decreases. Denoting the transmit power at the $i$-th AO iteration by $P\left(\boldsymbol{S}[i],\boldsymbol{\theta}[i]\right)$. We have
[…]
where $(a)$ holds since the transmit power has a lower bound, $(b)$ holds since $\boldsymbol{S}[i+1]$ is the optimal BS beamforming matrix of the $i+1$-th AO iteration; $(c)$ holds because the transmit power only depends on $\boldsymbol{S}$.
IV-D Convergence Analysis for the AO-GEMM Algorithm
In this subsection, we analyze the convergence performance of the proposed AO-GEMM algorithm, i.e., the AO based on the SDR method and the RIS beamforming scheme 2. Since the analysis is similar with that of the last subsection, we omit the convergence analysis of the AO-GEMM algorithm.
IV-E Discussion
The complexity of the SDR method for solving the transmit beamforming scales approximately with the fourth power of the number of BS antennas, communication users, and sensing targets. When the numbers of BS antennas, users, and targets is large, considerable computational resources are required to execute the SDR method. In cases where computational resources are insufficient, another alternative suboptimal algorithm is needed to obtain the transmit beamforming matrix. Moreover, the complexity of the SDR method for solving the RIS beamforming scales approximately with the fourth power of the number of RIS elements and communication users. In cases where computational resources are insufficient, we recommend using the proposed GEMM algorithm, i.e., RIS beamforming scheme 2, to solve the RIS beamforming problem.
@@ V Simulation Results
We assume that the BS is located at(0m, 0m); the RIS is located at (50m,10m); the $K$ users and $L$ targets are uniformly distributed within a circle centered at (70m,0m) with a radius of 5 meters and within a circle centered at the base station with a radius of 70 meters. Unless otherwise specified, the number of ISAC BS antennas is 16, the number of targets is 2, the number of users is 2, the communication rate threshold $R_{0}$ equals 2 bits/s/Hz, the CRB threshold equals 0.01, the probability $\rho_{k}$ and $p_{l}$ both equal 0.05, and the noise power is -110dBm. Moreover, from Fig. 2 to Fig. 7, the transmit power obtained based on the AO-SDR algorithm and continuous RIS phase shifts is taken as the lower bound, labelled as “Continuous RIS phase shifts” because the AO-SDR algorithm has been shown in [11] to achieve a near-optimal solution.
Figure 2: Feasibility rate of the proposed AO algorithm versus the number of BS antennas when $\gamma_{\text{BU},k},\gamma_{\text{BRU},k}$, and $\varepsilon_{l}$ both equal 0.01.
Figure 3: Transmit power versus the number of BS antennas when $\gamma_{\text{BU},k},\gamma_{\text{BRU},k}$, and $\varepsilon_{l}$ both equal 0.01.
Figure 2 illustrates that the feasibility of all schemes is always high. This is because the BS active beamforming provides higher DoF to compensate for CSI errors, especially when the number of BS antennas is large.
Figure 3 shows that the required transmit power decreases as the number of BS antennas increases. This is because the beamforming gain provided by the BS antennas increases as the number of antennas increases, exceeding the power loss caused by compensating for CSI errors.
Figure 4 illustrates that all schemes achieve 100% feasibility when the number of RIS elements is below 34. However, as the number of RIS elements increases beyond 34, the feasibility of all schemes rapidly decreases. Only the channel realizations with good channel conditions are feasible when the number of RIS elements is large.
@@
The trends of all curves in Fig. 6 are consistent. When the number of RIS elements is small (less than about 10), the transmit power decreases as the number of RIS elements increases. However, when the number of RIS elements is larger than about 10, the transmit power increases with the number of RIS elements. This is because when the number of RIS elements is small, the transmit power used to compensate for CSI error is also small. Nevertheless, when the number of RIS elements is large (less than about 50), the power loss caused by compensating for CSI error is greater than the power gain provided by the RIS, resulting in an increase of the transmit power as the number of RIS elements increases. When the number of RIS elements is relatively large (greater than about 50), the feasibility probability of the proposed algorithm is low and only channel realizations with good channel conditions are feasible. In such channel realizations, the transmit power required for robust transmission is low. Therefore, the transmission power gradually decreases as the number of RIS elements increases beyond 50. These trends in Fig. 6 break the intuitive conclusion that the more RIS elements, the greater the gain. When perfect CSI cannot be obtained, we should carefully adjust the number of RIS elements. Moreover, the transmit power obtained by RIS beamforming scheme 1 is slightly higher than the lower bound because the discrete phase shift projection introduces quantization errors, which degrade system performance. The transmit power obtained by RIS beamforming scheme 2 is higher than that obtained by RIS beamforming scheme 1 because RIS beamforming scheme 2 is a suboptimal algorithm with low computational complexity that iteratively solves the non-convex problem in (40a) using the gradient ascent method.
@@
Figure 10: The cumulative distribution function (CDF) of the number of iterations.The number of RIS elements and BS antennas both equal to 16, the number of RIS phase bits $d$ equals 4, $\gamma_{\text{BU},k},\gamma_{\text{BRU},k}$, and $\varepsilon_{l}$ both equal 0.01.
Figure 9 and Figure 10 depict the performance of the AO algorithms based on RIS beamforming scheme 1 and RIS beamforming scheme 2. All schemes achieve convergence within 20 AO iterations. Notably, the AO algorithm based on RIS beamforming scheme 1 achieves convergence in 20 iterations, while the AO algorithm based on RIS beamforming scheme 2 achieves convergence in only four iterations. RIS beamforming scheme 2 converges faster than RIS beamforming scheme 1 because the Gaussian randomization in the later scheme may lead to a degradation in system performance. Using a penalty function, RIS beamforming scheme 2 achieves faster convergence by directly iterating along the gradient direction within the feasible space.
@@ VI Conclusion
This paper presents a novel approach for designing robust transmission schemes in RIS-assisted ISAC systems. A decomposition-based large deviation inequality approach is proposed to approximate non-convex outage probability constraints for communication and non-convex probability constraints for the sensing CRB. The AO algorithm is then employed to solve the joint beamforming problem. Specifically, the transmit beamforming problem is formulated as a SDP problem. We introduce two schemes to solve the discrete RIS beamforming problem, which are the SDR-based algorithm with high complexity and high precision and the SP-based GEMM algorithm with low complexity and low precision. Simulation results demonstrate the excellent performance of the proposed AO algorithms. It should be noted that increasing the number of RIS elements does not lead to better performance when the perfect CSI is not available. Thus, we should carefully select the number of RIS elements to guarantee communication rate and CRB for sensing parameter estimation.
@@
The second part of (17) is rewritten as
@@
where we user the property of the trace-operator and the vec-operator [49, Eq. (521)], i.e., $\boldsymbol{a}^{T}\boldsymbol{Bc}=\text{Tr}(\boldsymbol{ac}^{T}\boldsymbol{B})=\text{vec}(\boldsymbol{ca}^{T})^{T}\text{vec}(\boldsymbol{B})$. We can formulate (59) as $2\text{Re}(\boldsymbol{t}_{k}^{H}\boldsymbol{e}_{k})$ using the definition in (19) and (21).
The third part of (17) is rewritten as
@@
where we use the property of the vec-operator [49, Eq. (524)], i.e., $\boldsymbol{a}^{T}\boldsymbol{XBX}^{T}\boldsymbol{c}=\text{vec}(\boldsymbol{X})^{T}\left(\boldsymbol{B}\otimes\boldsymbol{ca}^{T}\right)\text{vec}(\boldsymbol{X})$. Please refer to B.1.1 of [49] for a proof for this property. We can formulate (60) as $\boldsymbol{e}_{k}^{H}\boldsymbol{Q}_{k}\boldsymbol{e}_{k}$ using the definition in (19) and (20).
The first part and the last part of (17) form $s_{k}$. Thus, we obtain (18). This completes the proof of Proposition 1.
@@
In order to derive a convex restriction of (18) using Lemma 1, we set $\boldsymbol{x}=\boldsymbol{e}_{k}$, $\boldsymbol{Q}=\boldsymbol{Q}_{k}$, $\boldsymbol{r}=\boldsymbol{r}_{k}$, $\eta_{k}=\text{Tr}(\boldsymbol{Q}_{k})+s_{k}$, and $T_{k}=v_{k}\|\boldsymbol{Q}_{k}\|_{F}+\frac{1}{\sqrt{2}}\|\boldsymbol{r}_{k}\|$. We choose $v_{k}$ as the solution to $[1-1/(2v_{k}^{2})]v_{k}=\sqrt{\ln(1/\rho_{k})}$. We could observe that when $\eta_{k}\textgreater 2\bar{\theta}_{k}v_{k}T_{k}$, i.e., $\eta_{k}\geq 2\sqrt{\ln(1/\rho_{k})}T_{k}$, the outage probability constraint (18) is always satisfied. It is shown as the following equation.
@@
Further, (18) is also satisfied if $\eta_{k}=2\bar{\theta}_{k}v_{k}T_{k}$. However, (18) may not be satisfied if $\eta_{k}\textless 2\bar{\theta}_{k}v_{k}T_{k}$. Thus, we add constraint for this situation to ensure the outage probability is no more that $\rho_{k}$. The constraint is expressed as